Bexley:

How many triangles can be drawn with one 90-angle, 70-angle, and an included side measuring 3 inches? Explain.

Aries:

You can only draw 2. If leg lengths are 5 and 12 then the only such right triangle is the famous 5-12-13 Pythagorean Triple. If 12 is the hypotenuse and a leg has a length of 5, then there's another right triangle. Unknown leg length x would satisfy 12^2 = 5^2 + x^2. IT LOOKS AS IF THERE ARE TWO possible such right triangles.

Bexley:

@aries wrote:
You can only draw 2. If leg lengths are 5 and 12 then the only such right triangle is the famous 5-12-13 Pythagorean Triple. If 12 is the hypotenuse and a leg has a length of 5, then there's another right triangle. Unknown leg length x would satisfy 12^2 = 5^2 + x^2. IT LOOKS AS IF THERE ARE TWO possible such right triangles.
thx lol

Aries:

Np